# The lattice structure of negative Sobolev and extrapolation spaces

### v1

@@ Abstract.
It is well-known that the Sobolev spaces $W^{k,p}(\mathbb{R}^{d})$ are vector lattices with respect to the pointwise almost everywhere order if $k\in\{0,1\}$, but not if $k\geq 2$. In this note, we consider negative $k$ and show that the span of the positive cone in $W^{k,p}(\mathbb{R}^{d})$ is a vector lattice in this case.
We also prove a related abstract result: if $(T(t))_{t\in[0,\infty)}$ is a positive $C_{0}$-semigroup on a Banach lattice $X$ with order continuous norm, then the span of the cone $X_{-1,+}$ in the extrapolation space $X_{-1}$ is a vector lattice. This complements results obtained by Bátkai, Jacob, Wintermayr, and Voigt in the context of perturbation theory and provides additional context for the theory of infinite-dimensional positive systems.
@@ 1. Introduction
For every $p\in[1,\infty]$, the first order Sobolev space $W^{1,p}(\mathbb{R}^{d})$ is a vector lattice with respect to the pointwise almost everywhere order. This follows from the fact that $\left\lvert f\right\rvert\in W^{1,p}(\mathbb{R}^{d})$ for all $f\in W^{1,p}(\mathbb{R}^{d})$ [9, Lemma 7.6]; in other words, the Sobolev space is even a sublattice of the Banach lattice $L^{p}(\mathbb{R}^{d})$. In contrast, $W^{k,p}(\mathbb{R}^{d})$ is not a vector lattice for $k\geq 2$ [2, Example 2.3(d)]; however, the cone is generating. In this note we are interested in the case of negative $k$. At first glance, the situation seems to be trivial: for an integer $k<0$, there are also elements of $W^{k,p}(\mathbb{R}^{d})$ which cannot be written as the difference of two positive elements. In particular, the cone is not generating, whence the space is not a vector lattice.
@@
However this simple answer does not really give us much insight into the order structure of $W^{k,p}(\mathbb{R}^{d})$ for negative $k$ – it simply means that the cone is small compared to the entire space (which is, from a theoretical point of view, a consequence of the fact that the cone in Sobolev spaces of positive order is not normal). If we restrict our attention to the span $W^{k,p}(\mathbb{R}^{d})_{+}-W^{k,p}(\mathbb{R}^{d})_{+}$ of the positive cone, we shall see in Section 3 that this is indeed a vector lattice for every integer $k<0$.
@@
A related phenomenon occurs, in a more abstract setting, in the theory of positive operator semigroups. If $X$ is a Banach lattice and $T=(T(t))_{t\geq 0}$ is a positive $C_{0}$-semigroup on $X$, one can associate a so-called extrapolation space $X_{-1}$ to $T$ – this is a larger space into which $X$ embeds continuously and densely. The space $X_{-1}$ can be considered as an abstract Sobolev space of negative order. To get an order structure on $X_{-1}$ it is natural to define the cone $X_{-1,+}$ as the closure of $X_{+}$ within the space $X_{-1}$. This order structure was introduced by Bátkai, Jacob, Wintermayr, and Voigt in [4] in order to study positive unbounded perturbations of semigroup generators; they also proved that this is indeed a cone and that $X_{+}=X_{-1,+}\cap X$, i.e., the order structures on $X$ and $X_{-1}$ are compatible [4, Proposition 2.3]. The same order on $X_{-1}$ plays a significant role in the theory of positive linear systems in infinite-dimensions: there, one is interested in positive control operators $B$ that operate from an ordered Banach space $U$, the so-called input space, into the extrapolation space $X_{-1}$; see the recent papers [7, 3] and also [8]. Just as for the concrete Sobolev space $W^{k,p}(\mathbb{R}^{d})$ for $k<0$, the cone $X_{-1,+}$ does not span $X_{-1}$, in general, and hence $X_{-1}$ is not a vector lattice. This was observed in a concrete example in [4, Example 5.1]. Yet, we show in Section 4 that a similar phenomenon as for the concrete Sobolev spaces occurs, i.e., the span $X_{-1,+}-X_{-1,+}$ is indeed a vector lattice provided that the Banach lattice $X$ has order continuous norm.
@@
Our results are based on general results on how to transfer lattice properties between ordered Banach spaces that are not isomorphic. We prove those results in Section 2. Different results regarding when an ordered Banach space (or ordered vector space) is automatically a vector lattice can be found in [10, Section 4].
@@
We close this introduction by recalling of few notions from the theory of ordered Banach spaces and Banach lattices.
@@
By an ordered Banach space, we mean a Banach space $X$ which is also an ordered vector space whose positive cone $X_{+}$ is closed with respect to the norm topology. The cone of an ordered Banach space is called normal if there exists a number $M\geq 1$ such that $0\leq x\leq y$ implies $\left\lVert x\right\rVert\leq M\left\lVert y\right\rVert$ for all $x,y\in X$. If $X$ is an ordered Banach space, then it is easy to see that $\operatorname{span}X_{+}=X_{+}-X_{+}$. Endowing this space with the norm
$\left\lVert x\right\rVert_{\operatorname{span}X_{+}}:=\inf\{\left\lVert y\right\rVert+\left\lVert z\right\rVert:\,y,z\in X_{+}\text{ and }x=y-z\}$ (1.1)
makes $\operatorname{span}X_{+}$ an ordered Banach space and makes the embedding $\operatorname{span}X_{+}\hookrightarrow X$ continuous [2, Lemma 2.2]. Moreover, on the cone $X_{+}$ this norm coincides with the norm on $X$. Hence, if the cone $X_{+}$ is normal in $X$, then it is also normal with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X_{+}}$.
An ordered vector space $X$ is called a vector lattice if any two elements have a supremum (equivalently, an infimum). The positive cone of a vector lattice is always generating and each element $x\in X$ has a modulus that is defined as $\left\lvert x\right\rvert:=\sup\{-x,x\}$. An ordered Banach space is called a Banach lattice if it is a vector lattice and the order is compatible with the norm in the following sense that
@@
for all $x,y\in X$. The cone of a Banach lattice is always normal. If $X$ is an ordered Banach space and $\operatorname{span}X_{+}$ is norm dense in $X$, then the so called dual wedge
@@
is also a cone and thus turns the dual space $X^{\prime}$ into an ordered Banach space (if $\operatorname{span}X_{+}$ is not dense in $X$, then $X^{\prime}_{+}\cap-X^{\prime}_{+}$ will may not be zero, so $X^{\prime}$ is then only a pre-ordered Banach space). If $X$ is a Banach lattice, then the dual space is a Banach lattice as well.
For ordered Banach spaces $X$ and $Z$ we denote the space of all bounded linear operators from $X$ to $Z$ by $\mathcal{L}(X,Z)$ and we abbreviate $\mathcal{L}(X):=\mathcal{L}(X,Z)$. A linear map $T:X\to Z$ is called positive if $TX_{+}\subseteq Z_{+}$. As a consequence of the Hahn-Banach separation theorem, this is equivalent to $T^{\prime}Z^{\prime}_{+}\subseteq X^{\prime}_{+}$. By $\mathcal{L}(X,Z)_{+}$ we denote the set of all positive operators in $\mathcal{L}(X,Z)$. The identity operator on $X$ will be denoted by $\operatorname{id}_{X}$.
@@ 2. Abstract results on ordered Banach spaces
Recall that a Banach lattice $Z$ is called monotonically complete if every increasing norm-bounded net in $Z$ (equivalently, in $Z_{+}$) has a supremum [11, Definition 2.4.18]. Every $L^{p}$-space is monotonically complete for $p\in[1,\infty)$ and $L^{\infty}$-spaces over a $\sigma$-finite measure space are also monotonically complete. Clearly, a monotonically complete Banach lattice is Dedekind complete, i.e., every order-bounded increasing net in such a space has a supremum. The space $c_{0}$ of sequences that converge to 00 shows that the converse implication does not hold. A KB-space is a Banach lattice in which every increasing norm-bounded net (equivalently, sequence) is norm-convergent. Typical examples of KB-spaces are all reflexive Banach lattices and $L^{1}$-spaces (i.e., all $L^{p}$-spaces for $1\leq p<\infty$).
@@
Let $X$ be an ordered Banach space with generating cone and let $Z$ be a Banach lattice. Assume that there exists an operator $J\in\mathcal{L}(X,Z)_{+}$ and a sequence $(R_{n})\subseteq\mathcal{L}(Z,X)_{+}$ such that $(R_{n}J)\subseteq\mathcal{L}(X)$ converges to the $\operatorname{id}_{X}$ with respect to the weak operator topology.
Then $\operatorname{span}(X^{\prime}_{+})=X^{\prime}_{+}-X^{\prime}_{+}$ is a monotonically complete Banach lattice with respect to a norm equivalent $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X^{\prime}_{+}}$. In addition, if $X$ is reflexive then the Banach lattice $\operatorname{span}(X^{\prime}_{+})$ is a KB-space.
We point out that, in case of $X$ being reflexive, the conclusion that $\operatorname{span}X^{\prime}_{+}$ is a KB-space cannot be strengthened to $\operatorname{span}X^{\prime}_{+}$ also being reflexive; see the discussion after Corollary 2.4 below. Note that, if a sequence in $X_{+}$ converges in $X$, then it may not necessarily converge in $\operatorname{span}X_{+}$ as well. However, this is the case for increasing sequences. We state this observation in the following lemma as it is used multiple times in the sequel.
@@
Let $X$ be an ordered Banach space and let $(x_{n})$ be an increasing net of elements in $\operatorname{span}X$. If $(x_{n})$ is norm-convergent in $X$, then it is also norm convergent with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}X_{+}}$.
@@
Let $(x_{n})$ be norm convergent in $X$, to say $x$. Since $(x_{n})$ is increasing, $x-x_{n}\in X_{+}$ for each index $n$. As the norm on $X$ and the norm on $\operatorname{span}X_{+}$ coincide on $X_{+}$, we obtain that
$\left\lVert x_{n}-x\right\rVert_{\operatorname{span}X_{+}}=\left\lVert x_{n}-x\right\rVert_{X}\to 0$
as $n\to\infty$. ∎
@@
Step 1: The span is a lattice. Let $x^{\prime}\in\operatorname{span}X^{\prime}_{+}$. It suffices to prove that ±$\pm x^{\prime}$ has a supremum in $\operatorname{span}X^{\prime}_{+}$. Since $Z$ is a Banach lattice, so is its dual space $Z^{\prime}$. Hence, we can take the modulus $\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert$ in $Z^{\prime}$ for each index $n$. We show that the sequence $\big{(}J^{\prime}\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert\big{)}$ is norm-bounded in $X^{\prime}$. First of all, because $(R_{n}J)$ converges weakly to $\operatorname{id}_{X}$, the sequence $(J^{\prime}R_{n}^{\prime})$ converges to $\operatorname{id}_{X^{\prime}}$ in the weak-operator topology. In particular, it is bounded. Secondly, as $x^{\prime}\in X^{\prime}_{+}-X^{\prime}_{+}$, there exists a vector $y\in X^{\prime}_{+}$ such that ±$\pm x^{\prime}\leq y^{\prime}$. Therefore, ±$\pm R_{n}^{\prime}x^{\prime}\leq R_{n}^{\prime}y^{\prime}$ in $Z^{\prime}$ and hence, $J^{\prime}\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert\leq J^{\prime}R_{n}^{\prime}y^{\prime}$ in $X^{\prime}$ for each $n$.
@@
By the Banach-Alaoglu theorem, we can find a weak-convergent subsequence $\big{(}J^{\prime}\left\lvert R_{n_{j}}^{\prime}x^{\prime}\right\rvert\big{)}$ with limit $s^{\prime}$ in $X^{\prime}$. As the cone $X^{\prime}_{+}$ is weak-closed, it follows that $s^{\prime}\geq 0$. Let us show that $s^{\prime}$ is the supremum of ±$\pm x^{\prime}$ in $\operatorname{span}X^{\prime}_{+}$: for every upper bound $u^{\prime}$ of ±$\pm x^{\prime}$ in $\operatorname{span}X^{\prime}_{+}$ – repeating the above argument with $u^{\prime}$ instead of $y^{\prime}$ – we get
@@
in $X^{\prime}$ for each index $j$. Employing weak-convergence of $(J^{\prime}R_{n}^{\prime})$ and weak-closedness of $X^{\prime}_{+}$, it thus follows that $s^{\prime}\leq u^{\prime}$, so $s^{\prime}$ is indeed the least upper bound of ±$\pm x^{\prime}$ in $\operatorname{span}X^{\prime}_{+}$.

### v3

@@ Abstract.
It is well-known that the Sobolev spaces $W^{k,p}(\mathbb{R}^{d})$ are vector lattices with respect to the pointwise almost everywhere order if $k\in\{0,1\}$, but not if $k\geq 2$. In this note, we consider negative indices $-k$ and show that the span of the positive cone in $W^{-k,p}(\mathbb{R}^{d})$ is a vector lattice in this case. On bounded domains $\Omega\subseteq\mathbb{R}^{d}$ we obtain a partial result in this direction.
We also prove a related abstract theorem: if $(T(t))_{t\in[0,\infty)}$ is a positive $C_{0}$-semigroup on a Banach lattice $X$ with order continuous norm, then the span of the cone $X_{-1,+}$ in the extrapolation space $X_{-1}$ is a vector lattice. This complements results obtained by Bátkai, Jacob, Wintermayr, and Voigt in the context of perturbation theory and provides additional context for the theory of infinite-dimensional positive systems.
@@ 1. Introduction
For every $p\in[1,\infty]$, the first order Sobolev space $W^{1,p}(\mathbb{R}^{d})$ is a vector lattice with respect to the pointwise almost everywhere order. This follows from the fact that $\left\lvert f\right\rvert\in W^{1,p}(\mathbb{R}^{d})$ for all $f\in W^{1,p}(\mathbb{R}^{d})$; see [10, Lemma 7.6]. In other words, the Sobolev space is even a sublattice of the Banach lattice $L^{p}(\mathbb{R}^{d})$. In contrast, for $k\geq 2$, the space $W^{k,p}(\mathbb{R}^{d})$ is not a vector lattice [2, Example (d) on Page 419]. In this note, we are interested in the case of negative $k$. At first glance, the situation seems to be trivial: for an integer $k>0$, there are elements of $W^{-k,p}(\mathbb{R}^{d})$ which cannot be written as the difference of two positive elements. In particular, the space is not a vector lattice.
@@
However this simple answer does not really give us much insight into the order structure of $W^{-k,p}(\mathbb{R}^{d})$ – it simply means that the cone is small compared to the entire space (which is, from a theoretical point of view, a consequence of the fact that the cone in Sobolev spaces of positive order is not normal; see the end of the introduction for the definition of this and further notions). If we restrict our attention to the span $W^{-k,p}(\mathbb{R}^{d})_{+}-W^{-k,p}(\mathbb{R}^{d})_{+}$ of the positive cone, we shall see in Theorem 3.1 that this is indeed a vector lattice. On bounded open sets $\Omega\subseteq\mathbb{R}^{d}$ the situation is more subtle and we prove a partial result for this case in the rest of Section 3.
@@
A related phenomenon occurs, in a more abstract setting, in the theory of positive operator semigroups. If $X$ is a Banach lattice and $T=(T(t))_{t\geq 0}$ is a positive $C_{0}$-semigroup on $X$, one can associate a so-called extrapolation space $X_{-1}$ to $T$ – this is a larger space into which $X$ embeds continuously and densely. The space $X_{-1}$ can be considered as an abstract Sobolev space of negative order. To get an order structure on $X_{-1}$ it is natural to define the cone $X_{-1,+}$ as the closure of $X_{+}$ within the space $X_{-1}$. This order structure was introduced by Bátkai, Jacob, Wintermayr, and Voigt in [5] in order to study positive unbounded perturbations of semigroup generators; they also proved that this is indeed a cone and that $X_{+}=X_{-1,+}\cap X$, i.e., the order structures on $X$ and $X_{-1}$ are compatible [5, Proposition 2.3]. In addition to further results about positive perturbations [4], the order on $X_{-1}$ also plays a significant role in the theory of positive linear systems in infinite-dimensions: there one is interested in positive control operators $B$ that operate from an ordered Banach space $U$, the so-called input space, into the extrapolation space $X_{-1}$; see the recent papers [8, 3] and also [9]. Just as for the concrete Sobolev space $W^{k,p}(\mathbb{R}^{d})$ for $k<0$, the cone $X_{-1,+}$ does not span $X_{-1}$, in general, and hence $X_{-1}$ is not a vector lattice. This was observed in a concrete example in [5, Example 5.1]. Yet, we show in Section 4 that a similar phenomenon as for the concrete Sobolev spaces occurs, i.e., the span $X_{-1,+}-X_{-1,+}$ is indeed a vector lattice provided that the Banach lattice $X$ has order continuous norm.
@@
Our results are based on more abstract theorems about how to transfer lattice properties between ordered Banach spaces that are not isomorphic. We prove those abstract theorems in Section 2. Different results regarding when an ordered Banach space (or ordered vector space) is automatically a vector lattice can be found in [11, Section 4].
@@
We close this introduction by recalling few notions from the theory of ordered Banach spaces and Banach lattices.
@@
By an ordered Banach space, we mean a Banach space $X$ which is also an ordered vector space whose positive cone $X_{+}$ is closed with respect to the norm topology. The cone of an ordered Banach space is called normal if there exists a number $M\geq 1$ such that $0\leq x\leq y$ implies $\left\lVert x\right\rVert\leq M\left\lVert y\right\rVert$ for all $x,y\in X$. If $X$ is an ordered Banach space, then it is easy to see that $\operatorname{span}(X_{+})=X_{+}-X_{+}$. Endowing this space with the norm
$\left\lVert x\right\rVert_{\operatorname{span}(X_{+})}:=\inf\{\left\lVert y\right\rVert+\left\lVert z\right\rVert:\,y,z\in X_{+}\text{ and }x=y-z\}$ (1.1)
makes $\operatorname{span}(X_{+})$ an ordered Banach space and makes the embedding $\operatorname{span}(X_{+})\hookrightarrow X$ continuous [2, Lemma 2.2]. Moreover, on the cone $X_{+}$ this norm coincides with the norm on $X$. Hence, if the cone $X_{+}$ is normal in $X$, then it is also normal with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(X_{+})}$.
An ordered vector space $X$ is called a vector lattice if any two elements have a supremum (equivalently, an infimum). The positive cone of a vector lattice is always generating and each element $x\in X$ has a modulus that is defined as $\left\lvert x\right\rvert:=\sup\{-x,x\}$. An ordered Banach space is called a Banach lattice if it is a vector lattice and the order is compatible with the norm in the sense that
@@
for all $x,y\in X$. The cone of a Banach lattice is always normal. Moreover, the elements $x^{+}:=\sup\{x,0\}$ and $x^{-}:=\sup\{-x,0\}$ satisfy $x=x^{+}-x^{-}$, so the cone of a Banach lattice is also generating. If $X$ is an ordered Banach space and $\operatorname{span}(X_{+})$ is norm dense in $X$, then the so called dual wedge
@@
is also a cone and thus turns the dual space $X^{\prime}$ into an ordered Banach space (if $\operatorname{span}(X_{+})$ is not dense in $X$, then $X^{\prime}_{+}\cap-X^{\prime}_{+}$ will not be zero, so $X^{\prime}$ is then only a pre-ordered Banach space). If $X$ is a Banach lattice, then the dual space is a Banach lattice as well.
A linear map $T:X\to Z$ between two ordered vector spaces $X$ and $Z$ is called bipositive if, for all $x\in X$, one has $Tx\geq 0$ if and only if $x\geq 0$. For ordered Banach spaces $X$ and $Z$ we denote the space of all bounded linear operators from $X$ to $Z$ by $\mathcal{L}(X,Z)$ and we abbreviate $\mathcal{L}(X):=\mathcal{L}(X,X)$. A linear map $T:X\to Z$ is called positive if $TX_{+}\subseteq Z_{+}$. As a consequence of the Hahn-Banach separation theorem, this is equivalent to $T^{\prime}Z^{\prime}_{+}\subseteq X^{\prime}_{+}$. By $\mathcal{L}(X,Z)_{+}$ we denote the set of all positive operators in $\mathcal{L}(X,Z)$. The identity operator on $X$ will be denoted by $\operatorname{id}_{X}$.
@@ 2. Abstract results on ordered Banach spaces
Recall that a Banach lattice $Z$ is called monotonically complete if every increasing norm-bounded net in $Z$ (equivalently, in $Z_{+}$) has a supremum [14, Definition 2.4.18]. Every $L^{p}$-space is monotonically complete for $p\in[1,\infty)$ and $L^{\infty}$-spaces over a $\sigma$-finite measure space are also monotonically complete. Clearly, a monotonically complete Banach lattice is Dedekind complete, i.e., every order-bounded increasing net in such a space has a supremum. The space $c_{0}$ of sequences that converge to 00 shows that the converse implication does not hold. A KB-space is a Banach lattice in which every increasing norm-bounded net (equivalently, sequence) is norm-convergent. Typical examples of KB-spaces are all reflexive Banach lattices and $L^{1}$-spaces (i.e., all $L^{p}$-spaces for $1\leq p<\infty$). A subspace $I$ of a Banach lattice $Z$ is said to be lattice ideal if for each $x,y\in Z$, the conditions $0\leq x\leq y$ and $y\in I$ imply that $x\in I$.
@@
Let $X$ be an ordered Banach space with generating cone and let $Z$ be a Banach lattice. Assume that there exists an operator $J\in\mathcal{L}(X,Z)_{+}$ and a sequence $(R_{n})\subseteq\mathcal{L}(Z,X)_{+}$ such that $(R_{n}J)\subseteq\mathcal{L}(X)$ converges to $\operatorname{id}_{X}$ with respect to the weak operator topology.
[…]
    The space $\operatorname{span}(X^{\prime}_{+})=X^{\prime}_{+}-X^{\prime}_{+}$ is a monotonically complete Banach lattice with respect to a norm equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(X^{\prime}_{+})}$.
[…]
    If $X$ is reflexive, then the Banach lattice $\operatorname{span}(X^{\prime}_{+})$ is a KB-space.
[…]
    If the sequence $(JR_{n})\subseteq\mathcal{L}(Z)$ converges to $\operatorname{id}_{Z}$ with respect to the weak operator topology, then $J^{\prime}$ is a lattice homomorphism from $Z^{\prime}$ to $\operatorname{span}(X^{\prime}_{+})$ and $J^{\prime}Z^{\prime}$ is a lattice ideal in $\operatorname{span}(X^{\prime}_{+})$.
The proof of Theorem 2.1 requires some preparation that we outsource to Lemmata 2.2, 2.3, and 2.4. Note that, if a sequence in $X_{+}$ converges in $X$, then it may not necessarily converge with respect to the stronger norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(X_{+})}$ as well. However, this is the case for increasing sequences. This observation is the content of the following lemma.
@@
Let $X$ be an ordered Banach space and let $(x_{j})$ be an increasing net of elements in $\operatorname{span}(X_{+})$. If $(x_{j})$ is norm-convergent in $X$, then its limit is also in $\operatorname{span}(X_{+})$ and the net is even convergent with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\operatorname{span}(X_{+})}$.
@@
Let $(x_{j})$ be norm convergent in $X$, to say $x$. Since $(x_{j})$ is increasing, $x-x_{j}\in X_{+}$ for each index $j$. It thus follows from the assumption $x_{j}\in\operatorname{span}(X_{+})$ that $x\in\operatorname{span}(X_{+})$, too. As the norm on $X$ and the norm on $\operatorname{span}(X_{+})$ coincide on $X_{+}$, we obtain
$\left\lVert x-x_{j}\right\rVert_{\operatorname{span}(X_{+})}=\left\lVert x-x_{j}\right\rVert_{X}\to 0$
as claimed. ∎
As a second ingredient for the proof of Theorem 2.1 we use the following simple lemma about bipositive maps between vector lattices.
Lemma 2.3.
Let $X,Z$ be vector lattices and let $J:X\to Z$ be a bipositive linear map. If $JX_{+}$ is a face in $Z_{+}$, then $J$ is a lattice homomorphism and $JX$ is a lattice ideal in $Z$.
Proof.
Let $x\in X$. We have $0\leq\left\lvert Jx\right\rvert\leq J\left\lvert x\right\rvert$ due to the positivity of $J$. Therefore, using that $JX_{+}$ is a face in $Z_{+}$, we obtain a $y\in X_{+}$ such that $\left\lvert Jx\right\rvert=Jy$. Hence,
±$\pm Jx\leq Jy\leq J\left\lvert x\right\rvert.$
Since $J$ is bipositive this implies that ±$\pm x\leq y\leq\left\lvert x\right\rvert$, so actually $y=\left\lvert x\right\rvert$. In turn, $\left\lvert Jx\right\rvert=Jy=J\left\lvert x\right\rvert$. So $J$ is indeed a lattice homomorphism.
Consequently, $JX$ is a vector sublattice of $Z$ and by using again that $JX_{+}$ is a face in $Z_{+}$, we get that $JX$ is even a lattice ideal in $Z$. ∎
Finally, we also need the following lemma about renorming of ordered Banach spaces that are vector lattices. The lemma seems to be a folklore result that is known to experts in vector lattice theory and is used on various occasions in the literature, but it is hard to find a detailed proof of it. Since the proof is a bit less obvious than one might expect at first glance, we include the details.
Lemma 2.4.
Let $X$ be an ordered Banach space that is also a vector lattice (with respect to the same order) and assume that the cone $X_{+}$ is normal. Then there exists an equivalent norm on $X$ that renders $X$ a Banach lattice.
Proof.
Let $\left\lVert\mathord{\,\cdot\,}\right\rVert$ denote the given norm on $X$. Since $X_{+}$ is normal there exists, by the definition of normality, a real number $M\geq 1$ such that $\left\lVert x\right\rVert\leq M\left\lVert y\right\rVert$ for all $x,y\in X$ that satisfy $0\leq x\leq y$. Moreover, as $X$ is a vector lattice its cone is generating, so there exists a number $C\geq 1$ such that each $x\in X$ can be decomposed as $x=y-z$ for vectors $y,z\in X_{+}$ that satisfy $\left\lVert x\right\rVert,\left\lVert y\right\rVert\leq C\left\lVert x\right\rVert$; see [1, Theorem 2.37(1) and (3)]. We define
[…]
for each $x\in X$. The supremum is finite due to the normality of the cone. We now check that $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\mathrm{BL}}$ is a norm with the claimed properties.
The mapping $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\mathrm{BL}}$ is indeed a norm: Clearly, $\left\lVert 0\right\rVert_{\mathrm{BL}}=0$. If, conversely, $\left\lVert x\right\rVert_{\mathrm{BL}}=0$ for a vector $x\in X$, then it follows from $\left\lVert\left\lvert x\right\rvert\right\rVert\leq\left\lVert x\right\rVert_{\mathrm{BL}}=0$ that $\left\lvert x\right\rvert=0$, so $x=0$. For each $x\in X$ one easily checks that $\left\lVert\alpha x\right\rVert_{\mathrm{BL}}=\alpha\left\lVert x\right\rVert_{\mathrm{BL}}$ for all $\alpha\in[0,\infty)$ and that $\left\lVert-x\right\rVert_{\mathrm{BL}}=\left\lVert x\right\rVert_{\mathrm{BL}}$, so absolute homogeneity also holds. The triangle inequality can be checked by using the Riesz decomposition property, i.e., the equality $[0,x+y]=[0,x]+[0,y]$ for all $x,y\in X_{+}$, which holds in every vector lattice [17, Proposition II.1.6].
The norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\mathrm{BL}}$ is equivalent to $\left\lVert\mathord{\,\cdot\,}\right\rVert$: Let $x\in X$. We have $x=x^{+}-x^{-}$, so it follows from $0\leq x^{+}\leq\left\lvert x\right\rvert$ and $0\leq x^{-}\leq\left\lvert x\right\rvert$ that
$\left\lVert x\right\rVert\leq\left\lVert x^{+}\right\rVert+\left\lVert x^{-}\right\rVert\leq 2M\left\lVert\left\lvert x\right\rvert\right\rVert\leq 2M\left\lVert x\right\rVert_{\mathrm{BL}}.$
On the other hand, as above, we can find vectors $y,z\in X_{+}$ that satisfy $x=y-z$ and $\left\lVert y\right\rVert,\left\lVert z\right\rVert\leq C\left\lVert x\right\rVert$. One has $0\leq x^{+}\leq y$ and $0\leq x^{-}\leq z$, so for all $w\in X$ that satisfy $0\leq w\leq\left\lvert x\right\rvert=x^{+}+x^{-}$ we get
$\left\lVert w\right\rVert\leq M\left\lVert\left\lvert x\right\rvert\right\rVert\leq M\big{(}\left\lVert x^{+}\right\rVert+\left\lVert x^{-}\right\rVert\big{)}\leq M^{2}\big{(}\left\lVert y\right\rVert+\left\lVert z\right\rVert\big{)}\leq 2M^{2}C\left\lVert x\right\rVert.$
Hence, $\left\lVert x\right\rVert_{\mathrm{BL}}\leq 2M^{2}C\left\lVert x\right\rVert$.
The space $X$ is a Banach lattice with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\mathrm{BL}}$: Since the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\mathrm{BL}}$ is equivalent to the original norm, $X$ is still a Banach space with respect to $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\mathrm{BL}}$. If $x,y\in X$ satisfy $\left\lvert x\right\rvert\leq\left\lvert y\right\rvert$, then
[…]
and hence, $\left\lVert x\right\rVert_{\mathrm{BL}}\leq\left\lVert y\right\rVert_{\mathrm{BL}}$. ∎
Now we have all the ingredients available that we need in order to show Theorem 2.1.
@@
(a) We divide the proof into two steps.
Step 1: The span is a lattice. Let $x^{\prime}\in\operatorname{span}(X^{\prime}_{+})$. It suffices to prove that ±$\pm x^{\prime}$ has a supremum in $\operatorname{span}(X^{\prime}_{+})$. Since $Z$ is a Banach lattice, so is its dual space $Z^{\prime}$. Hence, we can take the modulus $\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert$ in $Z^{\prime}$ for each index $n$. We show that the sequence $\big{(}J^{\prime}\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert\big{)}$ is norm-bounded in $X^{\prime}$. First of all, because $(R_{n}J)$ converges to $\operatorname{id}_{X}$ in the weak operator topology, the sequence $(J^{\prime}R_{n}^{\prime})$ converges to $\operatorname{id}_{X^{\prime}}$ in the weak-operator topology. In particular, it is bounded. Secondly, as $x^{\prime}\in X^{\prime}_{+}-X^{\prime}_{+}$, there exists a vector $y\in X^{\prime}_{+}$ such that ±$\pm x^{\prime}\leq y^{\prime}$. Therefore, ±$\pm R_{n}^{\prime}x^{\prime}\leq R_{n}^{\prime}y^{\prime}$ in $Z^{\prime}$ and hence, $J^{\prime}\left\lvert R_{n}^{\prime}x^{\prime}\right\rvert\leq J^{\prime}R_{n}^{\prime}y^{\prime}$ in $X^{\prime}$ for each $n$.
@@
By the Banach-Alaoglu theorem, there is a weak-convergent subnet $\big{(}J^{\prime}\left\lvert R_{n_{j}}^{\prime}x^{\prime}\right\rvert\big{)}$ with limit $s^{\prime}$ in $X^{\prime}$. As the cone $X^{\prime}_{+}$ is weak-closed, it follows that $s^{\prime}\geq 0$. Let us show that $s^{\prime}$ is the supremum of ±$\pm x^{\prime}$ in $\operatorname{span}(X^{\prime}_{+})$. On one hand, it follows from the inequality ±$\pm J^{\prime}R_{n_{j}}^{\prime}x^{\prime}\leq J^{\prime}\left\lvert R_{n_{j}}^{\prime}x^{\prime}\right\rvert$ for each $j$, from the convergence of $(J^{\prime}R_{n}^{\prime})$ to $\operatorname{id}_{X^{\prime}}$ in the weak-operator topology, and from the weak-closedness of $X^{\prime}_{+}$, that ±$\pm x^{\prime}\leq s^{\prime}$. On the other hand, for every upper bound $u^{\prime}$ of ±$\pm x^{\prime}$ in $\operatorname{span}(X^{\prime}_{+})$ – repeating the above argument with $u^{\prime}$ instead of $y^{\prime}$ – we get
@@
in $X^{\prime}$ for each index $j$. Employing again the convergence of $(J^{\prime}R_{n}^{\prime})$ in the weak-operator topology and the weak-closedness of $X^{\prime}_{+}$, it thus follows that $s^{\prime}\leq u^{\prime}$. So $s^{\prime}$ is indeed the least upper bound of ±$\pm x^{\prime}$ in $\operatorname{span}(X^{\prime}_{+})$.